Solve.
12(x-4)-(x+2)=(x+2)(x-4)
They book has the answer of 7 and 6.
Can you tell me how they got this answer?
I came up with x = 13 and 208.

Answer by Edwin McCravy(20086)

12(x - 4) - (x + 2) = (x + 2)(x - 4)

   12x - 48 - x - 2 = x² - 4x + 2x - 8

           11x - 50 = x² - 2x - 8

                  0 = x² - 13x + 42

                  0 = (x - 7)(x - 6)

                       x - 7 = 0;   x - 6 = 0
                           
                           x = 7        x = 6

Edwin